# Extrinsic contribution to bosonic thermal Hall transport

### v1

@@ Abstract
Bosonic excitations like phonons and magnons dominate the low-temperature transport of magnetic insulators. Similar to electronic Hall responses, the thermal Hall effect (THE) of charge neutral bosons has been proposed as a powerful tool for probing topological properties of their wavefunctions. For example, the intrinsic contribution of the THE of a perfectly clean system is directly governed by the distribution of Berry curvature, and many experiments on topological magnon and phonon insulators have been interpreted in this way. However, disorder is inevitably present in any material and its contribution to the THE has remained poorly understood. Here we develop a rigorous kinetic theory of the extrinsic side-jump contribution to the THE of bosons. We show that the extrinsic THE can be of the same order as the intrinsic one but sensitively depends on the type of local imperfection. We study different types of impurities and show that a THE can even arise as a pure impurity-induced effect in a system with a vanishing intrinsic contribution. As a side product, we also generalize existing results for the electronic AHE to general types of impurities beyond the standard assumption of local potential scattering. We discuss the importance of our results for the correct interpretation of THE measurements.
@@ I Introduction
For charged particles, such as electrons, the origin of classical Hall transport from the Lorentz force has been known since the end of the 19th century. After the discovery of the quantum Hall effect (QHE) a century later Klitzing et al. (1980), its origin from the intrinsic Berry curvature of the electron’s wavefunctions was swiftly established Thouless et al. (1982). Understanding the nonquantized anomalous Hall effect (AHE) in metallic materials with broken time reversal (TR) symmetry has proven much more complicated, because it is both difficult experimentally and theoretically cumbersome to disentangle contributions from intrinsic Berry phase effects and extrinsic scattering contributions Nagaosa et al. . Charge neutral excitations in solids, for example phonons, magnons or exotic fractionalized spinons, can also exhibit Hall transport but now in response to gradients in temperature (or energy density Luttinger (1964)). A quantized thermal Hall effect (THE) has been put forward as a powerful probe of exotic fractionalized excitations from intrinsic topological order of Mott insulators Kitaev (2006); Kane and Fisher (1997), signatures of which have been arguably observed experimentally Kasahara et al. (2018); Paul et al. (2024). Unquantized THE in insulators can also originate from Berry phase effects of more mundane phonon and magnon excitations, as recently reported in many experiments on magnets Onose et al. (2010); Ideue et al. (2012); Hirschberger et al. (2015a, b); Watanabe et al. (2016); Doki et al. (2018); Hentrich et al. (2019); Akazawa et al. (2020); Zhang et al. (2021); Chen et al. (2022); Lefrançois et al. (2022); Bruin et al. (2022); Czajka et al. (2023); Sharma et al. (2024) – for a recent review see Zhang et al. (2024). While the theory of the intrinsic (Berry curvature) contribution to this bosonic THE has been firmly established over the last decade Matsumoto and Murakami (2011); Qin et al. (2011, ); Matsumoto et al. ; Shitade (2014); Zhang et al. (2020); Mangeolle et al. , much less is known about the extrinsic contributions. Concretely, scattering contributions from disorder inevitably present in any material are not rigorously understood and the subject of the present paper.
@@
×$\qquad\qquad\qquad\times\varepsilon^{2}\partial_{\varepsilon}n_{\rm B}\left(\varepsilon,T\right)\,\mathsf{\Omega}_{p_{x}p_{y}}(p),$
where $\Theta$ is the Heaviside step function, ${\rm Tr}_{+}$ is a sum over boson bands, $n_{\rm B}(\varepsilon,T)=1/(e^{\varepsilon/T}-1)$ is the Bose function, $\underline{{\sf K}}_{\rm d}(p)$ is the bosonic energy, and crucially $\mathsf{\Omega}_{p_{x}p_{y}}(p)$ is the bosons’ Berry curvature. This general formula for the intrinsic thermal Hall conductivity of neutral bosons was first derived rigorously using the Kubo-Streda formalism for phonons Qin et al. and magnons Matsumoto et al. . Its intuitive picture as a semiclassical current carried by particles obeying the semiclassical equations of motion Matsumoto and Murakami (2011) was recently provided rigorous support in Ref. Mangeolle et al. , where one of the authors took part in deriving Eq.(1) directly from the inhomogeneous kinetic equation. Except for the now well-known complications introduced by the energy magnetization Qin et al. (2011), this development is very similar to that of the intrinsic anomalous Hall effect (AHE) of electrons, which was first obtained from the Kubo formula Karplus and Luttinger before it was given a semiclassical picture using the semiclassical equations of motion of the electron wavepacket Sundaram and Niu , later derived rigorously within the kinetic equation using the Wigner-Weyl formalism Wickles and Belzig .
In both the (bosonic) thermal and (fermionic) electrical cases, the intrinsic contribution is but part of the total Hall conductivity. Other phenomena contribute, that are usually classified as skew-scattering Mori et al. (2014); Flebus and MacDonald (2022); Sun et al. (2022); Mangeolle et al. (2022a, b); Chatzichrysafis and Mook (2025) and side-jump mechanisms: for a review in the electronic case see Nagaosa et al. . The effect of disorder on the electronic AHE, and especially the side-jump effect, was originally computed in specific electron models using Kubo’s formula Adams and Blount ; Fivaz ; Nozières and Lewiner , but can also be derived in the language of semiclassical dynamics. This was shown much later, based on the wavepacket analysis Sinitsyn et al. (2005, 2006), and was also derived rigorously in a particular model within the Schwinger-Keldysh kinetic formalism König and Levchenko . At least in one particular case, a formal connection can be drawn between both approaches Sinitsyn et al. (2007). The general expectation is that the electronic side-jump effect is non-perturbative, in that it does not depend on the strength of disorder (or impurity density), when inter-particle interactions can be neglected. Yet, for electronic systems, the side-jump effect is actually not so commonly seen experimentally  [^1] See Nagaosa et al. and especially, in Sec. I.B.3: “A practical approach, which is followed at present for materials in which $\sigma^{\rm AH}$ seems to be independent of $\sigma_{xx}$, is to first calculate the intrinsic contribution to the AHE. If this explains the observation (and it appears that it usually does), then it is deemed that the intrinsic mechanism dominates. If not, we can take some comfort from understanding on the basis of simple model results that there can be other contributions to $\sigma^{\rm AH}$ which are also independent of $\sigma_{xx}$ and can for the most part be identified with the side-jump mechanism., because electron-electron interactions are usually strong especially in two dimensions Abrahams et al. (1981) and dominate over other scattering effects due e.g. to disorder.
In the case of the thermal Hall conductivity of neutral bosons, the side-jump effect was seldom studied (for a rare exception see Ref. Guo et al. (2022) within the linear response formalism Kapustin and Spodyneiko (2020)), and a general, intuitive, semiclassical picture based on kinetic theory is still missing. We note that a recent effort was made to include the effect of band broadening in the intrinsic thermal Hall effect Koyama and Nasu : this effect is perturbative in the disorder strength, and thus fully distinct from side-jump which appears as a separate, non-perturbative contribution. At odds with the electronic case, it is plausible that side-jump can play a significant role in the thermal Hall effect of phonons or magnons, insofar as interactions between bosons are usually small. For instance, they are due to lattice anharmonicity for phonons, or higher-order terms in the $1/S$ expansion for magnons, and they are usually local (as opposed to long-range like the Coulomb interaction between electrons).  [^2] In magnetic systems, long-range interactions also exist such as dipolar exchange, though instances where these dominate over other (local) magnetic interactions governing the physical properties are rare, especially for bulk 3D insulating magnets.
@@
Figure 1: Thermal Hall conductivity divided by temperature $\kappa_{xy}/T$ as a function of temperature $T$ for three different instances (a,b,c) of disorder and time reversal (TR) breaking parameters, in a minimal model of bosons with viscosity (TR-breaking) terms introduced in Sec. V. $\kappa_{xy}^{\rm clean}$ is the intrinsic contribution in the absence of disorder (Eq.(1)), $\kappa_{xy}^{\rm dis}$ is the contribution due to disorder (Eq.(2)), and $\kappa_{xy}^{\rm tot}=\kappa_{xy}^{\rm clean}+\kappa_{xy}^{\rm dis}$. Both axes are in arbitrary units, identical in all three plots. (a) Disorder is a local modulation of the clean Hamiltonian, with identical matrix structure. (b) Disorder is a local viscosity term, distinct from that already present in the clean theory. (c) The clean theory has zero viscosity and preserves TR, but impurities still introduce TR breaking locally: remarkably, in this case thermal Hall conductivity is a pure impurity effect.
@@
In this paper, we derive rigorously the side-jump effect in the thermal Hall conductivity of neutral bosons. Employing the kinetic equation formalism for systems of non-interacting bosons we include the presence of generic disorder fields, i.e. not only potential but a fully-fledged local deformation of the Hamiltonian as the most genereal form of disorder. Within the semiclassical approximations which we discuss later, we find that disorder scattering induces new contributions in the boson thermal Hall conductivity which are carried by bulk currents. As our main result we derive the corresponding disorder-induced thermal Hall conductivity reading
@@
×$\qquad\times p_{\lambda}\left[{\mathsf{\Omega}}^{\mathsf{W}}_{p_{y}p_{\lambda}}(p)\,\partial_{p_{x}}\underline{\mathsf{K}}_{\rm d}(p)-{\mathsf{\Omega}}^{\mathsf{W}}_{p_{x}p_{\lambda}}(p)\,\partial_{p_{y}}\underline{\mathsf{K}}_{\rm d}(p)\right],$
where ${\mathsf{\Omega}}^{\mathsf{W}}_{p_{\mu}p_{\mu}}(p)$ is a band-dependent and disorder-dependent quantity, generally distinct from the Berry curvature (and which reduces to it in specific cases). We find that the side-jump Hall effect is comparable in magnitude to the intrinsic Hall effect; because ${\mathsf{\Omega}}^{\mathsf{W}}_{p_{\mu}p_{\mu}}(p)$ depends on the instance of disorder considered, the above formula shows that $\kappa_{xy}^{\rm dis}$ is non-universal and depends on the specific instance of disorder that we consider, as is illustrated in Fig. 1. We highlight in particular case (c), which shows that even with vanishing intrinsic Berry curvature $\mathsf{\Omega}_{p_{x}p_{y}}$ there may still exist sizeable disorder-induced thermal Hall conductivity.
@@
Interestingly, our derivation translates with very little change to the case of electrical Hall transport by electrons. This provides a generalized formula for the side-jump AHE where impurities, as local random Hamiltonians, have a general matrix form in the space of fermion flavors and a general momentum dependence. We provide this electronic AHE contribution, together with a simple derivation, in Eq.(IV.5).
@@
The paper is organized as follows. In Section II, we summarize the derivation of Ref. Mangeolle et al. of the inhomogeneous kinetic theory without disorder, which forms the basis for the inclusion of disorder in a real-space, real-time picture. In Section III we expand the time evolution equation in phase space and show how disorder-induced coordinate shifts naturally appear in the full gauge-invariant kinetic equation. In Section IV we show how these coordinate shifts lead to a side-jump current operator, which we derive, and to a displaced equilibrium distribution, which together lead to Eq. (2). Finally, in Section V we apply our results to a concrete model of chiral bosons with different instances of disorder for illustration.

### v2

@@ Abstract
Bosonic excitations like phonons and magnons dominate the low-temperature transport of magnetic insulators. Similar to electronic Hall responses, the thermal Hall effect (THE) of charge neutral bosons has been proposed as a powerful tool for probing topological properties of their wavefunctions. For example, the intrinsic contribution of the THE of a perfectly clean system is directly governed by the distribution of Berry curvature, and many experiments on topological magnon and phonon insulators have been interpreted in this way. However, disorder is inevitably present in any material and its contribution to the THE has remained poorly understood. Here we develop a rigorous kinetic theory of the extrinsic side-jump contribution to the THE of bosons. We show that the extrinsic THE is always relevant for bosonic systems and can be of the same order as the intrinsic one but sensitively depends on the type of local imperfection. We study different types of impurities and show that a THE can even arise as a pure impurity-induced effect in a system with a vanishing intrinsic contribution. As a side product, we also generalize existing results for the electronic AHE to general types of impurities beyond the standard assumption of local potential scattering. We discuss the importance of our results for the correct interpretation of THE measurements, for example in the Kitaev magnet $\alpha$-RuCl3, and provide a ready-to-use formula for comparison to experimental data.
@@ I Introduction
For charged particles, such as electrons, the origin of classical Hall transport from the Lorentz force has been known since the end of the 19th century. After the discovery of the quantum Hall effect (QHE) a century later Klitzing et al. (1980), its origin from the intrinsic Berry curvature of the electron’s wavefunctions was swiftly established Thouless et al. (1982). Understanding the nonquantized anomalous Hall effect (AHE) in metallic materials with broken time reversal (TR) symmetry has proven much more complicated, because it is both difficult experimentally and theoretically cumbersome to disentangle contributions from intrinsic Berry phase effects and extrinsic scattering contributions Nagaosa et al. . Charge neutral excitations in solids, for example phonons, magnons or exotic fractionalized spinons, can also exhibit Hall transport but now in response to gradients in temperature (or energy density Luttinger (1964)). A quantized thermal Hall effect (THE) has been put forward as a powerful probe of exotic fractionalized excitations from intrinsic topological order of Mott insulators Kitaev (2006); Kane and Fisher (1997), signatures of which have been arguably observed experimentally Kasahara et al. (2018); Paul et al. (2024). Unquantized THE in insulators can also originate from more mundane phonon and magnon excitations, as recently reported in many experiments on magnets  Onose et al. (2010); Ideue et al. (2012); Hirschberger et al. (2015a, b); Watanabe et al. (2016); Doki et al. (2018); Hirokane et al. (2019); Akazawa et al. (2020); Zhang et al. (2021); Chen et al. (2022); Lefrançois et al. (2022), and particularly in the Kitaev material and spin liquid candidate $\alpha\text{-RuCl}_{3}$  Hentrich et al. (2019); Yamashita et al. (2020); Yokoi et al. (2021); Bruin et al. (2022); Kasahara et al. (2022); Zhang et al. (2023, 2024a); Czajka et al. (2023) — for a recent review see Zhang et al. (2024b). While the theory of the intrinsic (Berry curvature) contribution to this bosonic THE has been firmly established over the last decade Matsumoto and Murakami (2011); Qin et al. (2011, ); Matsumoto et al. ; Shitade (2014); Zhang et al. (2020); Mangeolle et al. , much less is known about the extrinsic contributions. Concretely, scattering contributions from disorder inevitably present in any material are still poorly understood and are the subject of the present paper.
@@
$\qquad\qquad\qquad\times\varepsilon^{2}\partial_{\varepsilon}n_{\rm B}\left(\varepsilon,T\right)\,\mathsf{\Omega}_{p_{x}p_{y}}(p),$
where $\Theta$ is the Heaviside step function, $\rm Tr$$_{+}$ is a sum over boson bands, $n_{\rm B}(\varepsilon,T)=1/(e^{\varepsilon/T}-1)$ is the Bose function, $\underline{{\sf K}}_{\rm d}(p)$ is the bosonic energy, and crucially $\mathsf{\Omega}_{p_{x}p_{y}}(p)$ is the bosons’ Berry curvature. This formula Eq.(1) is similar to the Karplus-Luttinger formula for the intrinsic AHE of electrons. However, in both the (bosonic) thermal and (fermionic) electrical cases, the intrinsic contribution is but part of the total Hall conductivity. Other phenomena contribute, that are usually classified as skew-scattering Mori et al. (2014); Flebus and MacDonald (2022); Sun et al. (2022); Mangeolle et al. (2022a, b); Chatzichrysafis and Mook (2025) and side-jump mechanisms: for a review in the electronic case see Nagaosa et al. . Here we investigate the latter.
@@
Figure 1: Thermal Hall conductivity divided by temperature $\kappa_{xy}/T$ as a function of temperature $T$ for three different instances (a,b,c) of disorder and time reversal (TR) breaking parameters, in a minimal model of bosons with viscosity (TR-breaking) terms introduced in Sec. V. $\kappa_{xy}^{\rm clean}$ is the intrinsic contribution in the absence of disorder (Eq.(1)), $\kappa_{xy}^{\rm dis}$ is the contribution due to disorder (Eq.(2)), and $\kappa_{xy}^{\rm tot}=\kappa_{xy}^{\rm clean}+\kappa_{xy}^{\rm dis}$. Both axes are in arbitrary units, identical in all three plots. (a) Disorder is a local modulation of the clean Hamiltonian, with identical matrix structure. (b) Disorder is a local viscosity term, distinct from that already present in the clean theory. (c) The clean theory has zero viscosity and preserves TR, but impurities still introduce TR breaking locally: remarkably, in this case thermal Hall conductivity is a pure impurity effect. (a,b,c) The temperature dependences and distinctive features of these curves are discussed in Sec.V.3.4.
The general expectation is that the electronic side-jump effect is non-perturbative, in that it does not depend on the strength of disorder (or impurity density), when inter-particle interactions can be neglected. Yet, for electronic systems, the side-jump effect is actually not so commonly seen experimentally  [^1] See Nagaosa et al. and especially, in Sec. I.B.3: “A practical approach, which is followed at present for materials in which $\sigma^{\rm AH}$ seems to be independent of $\sigma_{xx}$, is to first calculate the intrinsic contribution to the AHE. If this explains the observation (and it appears that it usually does), then it is deemed that the intrinsic mechanism dominates. If not, we can take some comfort from understanding on the basis of simple model results that there can be other contributions to $\sigma^{\rm AH}$ which are also independent of $\sigma_{xx}$ and can for the most part be identified with the side-jump mechanism., because electron-electron interactions are usually strong especially in two dimensions Abrahams et al. (1981) and dominate over other scattering effects due e.g. to disorder. In the presence of strong disorder, the electron liquid may also undergo localization in the bulk Abrahams et al. (1979), while mobile edge excitations account for a quantized Hall conductivity Prange and Girvin (1987): the side-jump contribution, as a bulk effect, is again negligible in this situation. Meanwhile, and at odds with the electronic case, the side-jump contribution is expected to play a significant role in the thermal Hall effect of phonons or magnons. Indeed, in the absence of Pauli exclusion, and even in the presence of a (mobility) gap and of edge states between two bulk bands, bosonic states from the lower bulk band are always thermally excited, so that they always contribute to thermal transport. Thus, side-jump effects are always present. Moreover, the latter are presumably much less dominated by interaction effects than in electron liquids, insofar as interactions between collective bosonic modes like phonons or magnons are usually small. For instance, they are due to lattice anharmonicity for phonons, or to higher-order terms in the $1/S$ expansion for magnons, and they are usually local (as opposed to long-range like the Coulomb interaction between electrons).  [^2] In magnetic systems, long-range interactions also exist such as dipolar exchange, though instances where these dominate over other (local) magnetic interactions governing the physical properties are rare, especially for bulk 3D insulating magnets. Possible experimental evidence for significant side-jump effects may be found in the kagome antiferromagnet Volborthite Watanabe et al. (2016), in pyrochlore terbium oxides Hirschberger et al. (2015a); Hirokane et al. (2019), and in the Kitaev compound $\alpha\text{-RuCl}_{3}$ Zhang et al. (2023) – see Sec. V.4.
@@
Motivated by many reports of THE in insulating materials Onose et al. (2010); Ideue et al. (2012); Hirschberger et al. (2015a, b); Watanabe et al. (2016); Doki et al. (2018); Hirokane et al. (2019); Akazawa et al. (2020); Zhang et al. (2021); Chen et al. (2022); Lefrançois et al. (2022); Hentrich et al. (2019); Yamashita et al. (2020); Yokoi et al. (2021); Bruin et al. (2022); Kasahara et al. (2022); Zhang et al. (2023, 2024a); Czajka et al. (2023); Zhang et al. (2024b), we derive rigorously the side-jump effect in the thermal Hall conductivity of neutral bosons. Employing a kinetic equation formalism (see below) for systems of non-interacting bosons, we include the presence of generic disorder fields, i.e. not only potential but a fully-fledged local deformation of the Hamiltonian as the most genereal form of disorder. Within the semiclassical approximations which we discuss later, we find that disorder scattering induces new contributions in the boson thermal Hall conductivity which are carried by bulk currents. As our main result we derive the corresponding disorder-induced thermal Hall conductivity reading
@@
$\qquad\times p_{\lambda}\left[{\mathsf{\Omega}}^{\mathsf{W}}_{p_{y}p_{\lambda}}(p)\,\partial_{p_{x}}\underline{\mathsf{K}}_{\rm d}(p)-{\mathsf{\Omega}}^{\mathsf{W}}_{p_{x}p_{\lambda}}(p)\,\partial_{p_{y}}\underline{\mathsf{K}}_{\rm d}(p)\right],$
@@
where ${\mathsf{\Omega}}^{\mathsf{W}}_{p_{\mu}p_{\mu}}(p)$ is a band-dependent and disorder-dependent quantity, generally distinct from the Berry curvature and which reduces to the latter in specific cases. Its expression is provided in Eq.(32), and a concrete numerical recipe how to compute it can be found in Appendix G. These show that ${\mathsf{\Omega}}^{\mathsf{W}}_{p_{\mu}p_{\mu}}$ depends on the detailed matrix structure of impurities: we discuss how the latter may be obtained for any given physical model in Sec.III.5. Interestingly, our derivation translates with very little change to the case of electrical Hall transport by electrons. This provides a generalized formula for the side-jump AHE where impurities, as local random Hamiltonians, have a general matrix form in the space of fermion flavors and a general momentum dependence. We provide this electronic AHE contribution, together with a simple derivation, in Eq.(IV.5).
As a concrete application, in Sec.V.2 we consider the honeycomb-lattice $K\Gamma\Gamma^{\prime}$ spin model in a magnetic field. This hamiltonian was proposed as a minimal model for the Kitaev material $\alpha\text{-RuCl}_{3}$ in Chern et al. (2021), where qualitative agreement was found with experimental measurements of $\kappa_{xy}$ in this compound. We investigate the role of disorder, in the form of random fluctuations of the model’s parameters that may be caused by chemical variations. We find that the side-jump term is a sizable contribution that leads to clear differences, both quantitative and qualitative, from the intrinsic THE, in accordance with experiments Czajka et al. (2023).
These observations appear to be quite generic. To illustrate this somewhat formally, in Sec.V.3 we investigate a generic low-energy theory for bosonic fields with broken time reversal symmetry. We find again that the side-jump thermal Hall effect is comparable in magnitude to the intrinsic thermal Hall effect (see Fig. 1). This makes Eq.(2) an indispensable contribution alongside Eq.(1). Since for bosonic systems the intrinsic and extrinsic contributions are not clearly seperated in different temperature regimes, i.e. they generically coexist, we argue that both necessarily need to be treated on equal footing for a correct interpretation of experimental data. Because ${\mathsf{\Omega}}^{\mathsf{W}}_{p_{\mu}p_{\mu}}(p)$ depends on the form of disorder considered, the above formula shows that $\kappa_{xy}^{\rm dis}$ is non-universal and depends on the specific instance of disorder that we consider, as illustrated in Fig. 1. We highlight in particular case (c), which shows that even with vanishing intrinsic Berry curvature $\mathsf{\Omega}_{p_{x}p_{y}}$ there may still exist sizeable disorder-induced thermal Hall conductivity.
While our paper discusses the physics of Eqs.(1),(2), both in generic and in concrete applications, a significant part of it is devoted to actually deriving Eq.(2) rigorously. Our proof uses the language of kinetic theory, thereby satisfying formal constraints (of systematicity and physical intuition) that we introduce next with some historical context. The general formula Eq.(1) for the intrinsic thermal Hall conductivity of neutral bosons was first derived rigorously using the Kubo-Streda formalism for phonons Qin et al. and magnons Matsumoto et al. . Its intuitive picture as a semiclassical current carried by particles obeying the semiclassical equations of motion Matsumoto and Murakami (2011) was recently provided rigorous support in Ref. Mangeolle et al. , where one of the authors took part in deriving Eq.(1) directly from the inhomogeneous kinetic equation. Except for the now well-known complications introduced by the energy magnetization Qin et al. (2011), this development is very similar to that of the intrinsic anomalous Hall effect (AHE) of electrons, which was first obtained from the Kubo formula Karplus and Luttinger before it was given a semiclassical picture using the semiclassical equations of motion of the electron wavepacket Sundaram and Niu , later derived rigorously within the kinetic language using the Wigner-Weyl formalism Wickles and Belzig . The effect of disorder on the electronic AHE, and especially the side-jump effect, was originally computed in specific electron models using Kubo’s formula Adams and Blount ; Fivaz ; Nozières and Lewiner , but can also be derived in the language of semiclassical dynamics. This was shown much later, based on the wavepacket analysis Sinitsyn et al. (2005, 2006), and was also derived rigorously in a particular model within the Schwinger-Keldysh kinetic formalism König and Levchenko . At least in one particular case, a formal connection can be drawn between both approaches Sinitsyn et al. (2007). In the case of the thermal Hall conductivity of neutral bosons, the side-jump effect was seldom studied (for a rare exception see Ref. Guo et al. (2022) within the linear response formalism Kapustin and Spodyneiko (2020)), and a general, intuitive, semiclassical picture based on kinetic theory was still missing – this is the problem we address in Sections II to IV.
@@
The rest of the paper is organized as follows. In Section II, we summarize the derivation of Ref.Mangeolle et al. of the inhomogeneous kinetic theory without disorder, which forms the basis for the inclusion of disorder in a real-space, real-time picture. We then show how disorder scattering can be included in this formalism. In Section III we expand the time evolution equation in phase space and show how disorder-induced coordinate shifts naturally appear in the full gauge-invariant kinetic equation. In Section IV we show how these coordinate shifts lead to a side-jump current operator, which we derive, and to a displaced equilibrium distribution, which together lead to Eq. (2). Finally, in Section V we apply our results, first very concretely to a spin model with direct experimental implications for the Kitaev compound $\alpha\text{-RuCl}_{3}$ Chern et al. (2021) (Sec.V.2), then to a more generic low-energy theory of chiral bosons (Sec.V.3), in both cases with different instances of disorder.